OpenStudy (anonymous):

the recommended dosage of the drug ampicillin for pediatric patients is 200 mg kilogram of the patients weight how many mg of ampicillin should be given to a patient who weighs 39 pounds

OpenStudy (nottim):

how much does a pediatric patient weigh?

OpenStudy (anonymous):

39 pounds

OpenStudy (nottim):

so then its ratios?

OpenStudy (nottim):

wait, converting mg into pounds?

OpenStudy (abb0t):

I thought 79 kg was avg weight for a male?

OpenStudy (campbell_st):

you need to change the 39 lb to kilograms.. 1kg is approx 2.205 lb so find weight in kg by dividing 39/2.205 =

OpenStudy (campbell_st):

once you know the weight in kilograms then its 200 mg x number ok kgs=

OpenStudy (whpalmer4):

Safest thing to do is include the units in the whole calculation. Dose = 200 mg per kg of body weight \[Dose = \frac{200 mg}{1 kg}* 39 lb * \frac{1 kg}{2.205 lb} = 3537.4 mg\] When all the units are canceled out, we're left with mg. If we set the equation up wrong, the units won't cancel in a sensible fashion.